OpenStudy (anonymous):

How many times does the 29th day of the month occur in 400 consecutive years?

OpenStudy (anonymous):

So it would be February 29 This is called a leap year. It occurs every four years.

OpenStudy (anonymous):

Yeah, I got that. So, Total number of months ( excluding February) + Februaries in leap years?

OpenStudy (anonymous):

Wouldn't the number of leap years in 400 years be 100? :?

OpenStudy (anonymous):

Every 4 years is a leap year... So: \[\frac{ 400 }{ 4}=100 \] \[100 \times 12=1200\]

OpenStudy (anonymous):

Yeah, so shouldn't it be 4500? But why does the answer say 4497?

OpenStudy (anonymous):

So then the leftover is: \[300 \times 11=3300\] Therefore: \[3300+1200= 4500\] So: 4500 times in 400 years

OpenStudy (zarkon):

you don't always have a leap year every 4 years

OpenStudy (anonymous):

really? huh didn't know that 0.o

OpenStudy (anonymous):

Oh, right. I got it. Like, for example : Year 100, 200 and 300 are not leap years?

OpenStudy (zarkon):

correct

OpenStudy (anonymous):

Oh right.... damn didn't see that lol thank you @Zarkon :)

OpenStudy (zarkon):

so during a 400 year period you should have 3 days that normally are leapyears but because that are divisible by 100 (and not 400) they are not

OpenStudy (anonymous):

When it comes to two 00s in the end, then you will divide it by 400 to check whether it is leap year or not..

OpenStudy (anonymous):

Like, in case of 1500, there is two 00s in then end so divide it by 400 here, not divisible so 1500 is not a leap year.. So do 100, 200, 300..

OpenStudy (anonymous):

Thank you, I think I get it now. ^^ So, it's 2500 - 3 = 4497.

OpenStudy (anonymous):

\[2500 - 3 = 2497\] lol error :)

OpenStudy (anonymous):

Lol, my bad. -.- I meant 4500 - 3. :P

OpenStudy (anonymous):

yup :P